OpenStudy (reneeshumpert):

A grocer wants to make a 10-pound mixture of peanuts and cashews that he can sell for $4.75 per pound. If peanuts cost $4.00 per pound and cashews cost $6.50 per pound, how many pounds of cashews should he use? 3 4 6 7 help?

OpenStudy (reneeshumpert):

you dont need to give me the answer Ive had this question before , just set me up alittle please

OpenStudy (luffingsails):

So the total cost of all of the nuts after they are mixed will be: Tcost = 10 * 4.75 = 47.50 It will be comprised of peanuts and cashews such that: 4.00 * (x nuts) + 6.50 * (y cashews) This mixture will equal the Total cost.

OpenStudy (luffingsails):

The other thing to think about is the relative cost of nuts and cashews. That's the final piece of the puzzle. Cashews = Nuts + 2.50. (or y = x+2.5).

OpenStudy (reneeshumpert):

so he would want 5 pounds of both , for a 10 pound mixture right ?

OpenStudy (reneeshumpert):

hold up nevermind. Im wrong

OpenStudy (luffingsails):

So, we have: Tcost = 47.50 4 * x + 6.5 * y = 47.5 and y = x + 2.5 So, solve: 4 * x + 6.5 * (x + 2.5) = 47.5

OpenStudy (luffingsails):

Simplifying: 4x + 6.5x + 16.25 = 47.5 10.5x = 31.25 x = 2.98 pounds. Then y = x + 2.5 y = 5.47 pounds.

OpenStudy (reneeshumpert):

okay, so what does y equal exactly you never explained. 5.75 pounds is an answer option. So does that represent the number of peanuts? therefore I would subtract that from 10 pounds to get my number of cashews?

OpenStudy (reneeshumpert):

isn't*

OpenStudy (luffingsails):

I'm sorry, Renee.. The other part of what we needed to know is exactly what you said (x + y = 10.0) So, instead of x+2.5 = y... use y = 10 - x... This gives instead.... 4x + 6.5(10-x) = 47.50

OpenStudy (luffingsails):

Sorry, I keep getting disconnected here at work. lol

OpenStudy (reneeshumpert):

so it would be 4x+ 65-6.5x=47.50 ?

OpenStudy (reneeshumpert):

65-2.5x=47.50 I mean

OpenStudy (luffingsails):

Right! x - 7 y = 10 - x = 10 - 7 = 3!

OpenStudy (reneeshumpert):

thank you !

OpenStudy (luffingsails):

You're welcome. Have a great morning.